# Multistep Consistency Models

### v1

@@ Abstract
In this paper we propose Multistep Consistency Models: A unification between Consistency Models (Song et al., 2023) and TRACT (Berthelot et al., 2023) that can interpolate between a consistency model and a diffusion model: a trade-off between sampling speed and sampling quality. Specifically, a 1-step consistency model is a conventional consistency model whereas we show that a $\infty$-step consistency model is a diffusion model.
Multistep Consistency Models work really well in practice. By increasing the sample budget from a single step to 2-8 steps, we can train models more easily that generate higher quality samples, while retaining much of the sampling speed benefits. Notable results are 1.4 FID on Imagenet 64 in 8 step and 2.1 FID on Imagenet128 in 8 steps with consistency distillation. We also show that our method scales to a text-to-image diffusion model, generating samples that are very close to the quality of the original model.
Machine Learning, ICML
@@ 2 Background: Diffusion Models
Diffusion models are specified by a destruction process that adds noise to destroy data: ${\bm{z}}_{t}=\alpha_{t}{\bm{x}}+\sigma_{t}{\bm{{\epsilon}}}_{t}$ where ${\bm{{\epsilon}}}_{t}\sim\mathcal{N}(0,1)$. Typically for $t\to 1$, ${\bm{z}}_{t}$ is approximately distributed as a standard normal and for $t\to 0$ it is approximately ${\bm{x}}$. In terms of distributions one can write the diffusion process as:
$q({\bm{z}}_{t}|{\bm{x}})=\mathcal{N}({\bm{z}}_{t}|\alpha_{t}{\bm{x}},\sigma_{t}).$ (1)
Following (Sohl-Dickstein et al., 2015; Ho et al., 2020) we will let $\sigma_{t}^{2}=1-\alpha_{t}^{2}$ (variance preserving). As shown in (Kingma et al., 2021), the specific values of $\sigma_{t}$ and $\alpha_{t}$ do not really matter. Whether the process is variance preserving or exploding or something else, they can always be re-parameterized into the other form. Instead, it is their ratio that matters and thus it can be helpful to define the signal-to-noise ratio, i.e. $\mathrm{SNR}(t)=\alpha_{t}^{2}/\sigma_{t}^{2}$.
To sample from these models, one uses the denoising equation:
$q({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{x}})=\mathcal{N}({\bm{z}}_{s}|\mu_{t\to s}({\bm{z}}_{t},{\bm{x}}),\sigma_{t\to s})$ (2)
@@
$||f({\bm{z}}_{t},t)-\operatorname{nograd}(f(\operatorname{DDIM}_{t\to s}({\bm{x}}_{\mathrm{teacher}},{\bm{z}}_{t}),s))||,$ (4)
@@
${\bm{z}}_{s}=\operatorname{DDIM}_{t\to s}({\bm{x}},{\bm{z}}_{t})=\alpha_{s}{\bm{x}}+(\sigma_{s}/\sigma_{t})({\bm{z}}_{t}-\alpha_{t}{\bm{x}})$ (5)
@@
${\bm{x}}=\operatorname{invDDIM}_{t\to s}({\bm{z}}_{s},{\bm{z}}_{t})=\frac{{\bm{z}}_{s}-\frac{\sigma_{s}}{\sigma_{t}}{\bm{z}}_{t}}{\alpha_{s}-\alpha_{t}\frac{\sigma_{s}}{\sigma_{t}}}.$ (6)
@@ 3 Multistep Consistency Models
Algorithm 1 Multistep Consistency Models
@@
Algorithm 2 Sampling from Multistep Consistency
  Sample ${\bm{z}}_{T}\sim\mathcal{N}(0,{\mathbf{I}})$
     ${\bm{z}}_{s}=\operatorname{DDIM}_{t\to s}(f({\bm{z}}_{t},t),{\bm{z}}_{t})$
  end for
@@ 3.1 General description
$||\operatorname{DDIM}_{t\to t_{\mathrm{step}}}(f({\bm{z}}_{t},t),{\bm{z}}_{t})-\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}}||,$ (7)
where $\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}}=\operatorname{DDIM}_{s\to t_{\mathrm{step}}}(\operatorname{nograd}f({\bm{z}}_{s},s))$ and where the teaching step ${\bm{z}}_{s}=\operatorname{aDDIM}_{t\to s}(x,{\bm{z}}_{t})$.
For now it suffices to think $\operatorname{aDDIM}$ as $\operatorname{DDIM}$. It will be described in detail in section 3.2. In fact, one can drop-in any deterministic sampler in place of $\operatorname{aDDIM}$ in the case of distillation.
@@
$||\hat{{\bm{x}}}_{\mathrm{diff}}||=||f({\bm{z}}_{t},t)-\operatorname{invDDIM}_{t\to t_{\mathrm{step}}}(\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}},{\bm{z}}_{t})||.$ (8)
@@
$||f({\bm{z}}_{t},t)-\operatorname{invDDIM}_{t_{\mathrm{step}}\to t}({\bm{z}}_{t_{\mathrm{step}}})||.$ (9)
@@
$\begin{split}w(t)||{\bm{x}}_{\mathrm{diff}}||&=\operatorname{invDDIM}_{t\to s}(\hat{{\bm{z}}}_{s},{\bm{z}}_{t})-\hat{{\bm{x}}}\\ &=w(t)||{\bm{x}}-\hat{{\bm{x}}}||\\ \end{split}$ (10)
@@
${\bm{z}}^{*}_{s}=\alpha_{s}{\bm{x}}^{*}+\sigma_{s}\frac{{\bm{z}}_{t}-\alpha_{t}{\bm{x}}^{*}}{\sigma_{t}}=(\alpha_{s}-\frac{\alpha_{t}\sigma_{s}}{\sigma_{t}}){\bm{x}}^{*}+\frac{\sigma_{s}}{\sigma_{t}}{\bm{z}}_{t}.$
@@
${\bm{z}}^{\text{DDIM}}_{s}=(\alpha_{s}-\alpha_{t}\sigma_{s}/\sigma_{t})\hat{{\bm{x}}}+(\sigma_{s}/\sigma_{t}){\bm{z}}_{t},$
@@
$||{\bm{z}}^{*}_{s}||^{2}-||{\bm{z}}^{\text{DDIM}}_{s}||^{2}=\text{trace}(\mathrm{Var}[{\bm{z}}_{s}|{\bm{z}}_{t}]),$
@@
$\mathrm{Var}[{\bm{z}}_{s}|{\bm{z}}_{t}]=(\alpha_{s}-\alpha_{t}\sigma_{s}/\sigma_{t})^{2}\cdot\mathrm{Var}[{\bm{x}}|{\bm{z}}_{t}],$
@@
Currently, the best sample quality is achieved with stochastic samplers, which can be tuned to add exactly enough noise to undo the oversmoothing caused by numerical integration. However, current distillation methods are not well suited to distilling these stochastic samplers directly. Here we therefore propose a new deterministic sampler that aims to achieve the norm increasing effect of noise addition in a deterministic way. It turns out we can do this by making a simple adjustment to the DDIM sampler, and we therefore call our new method Adjusted DDIM (aDDIM). Our modification is heuristic and is not more theoretically justified than the original DDIM sampler. However, empirically we find aDDIM to work very well leading to improved FID scores.
@@
${\bm{z}}^{\text{aDDIM}}_{s}=\alpha_{s}\hat{{\bm{x}}}+\sqrt{\sigma_{s}^{2}+\text{tr}(\mathrm{Var}[{\bm{z}}_{s}|{\bm{z}}_{t}])/||\hat{{\bm{{\epsilon}}}}||^{2}}\cdot\hat{{\bm{{\epsilon}}}}.$ (11)
@@ 4 Related Work
Existing works closest to ours are Consistency Models (Song et al., 2023; Song & Dhariwal, 2023) and TRACT (Berthelot et al., 2023). Compared to consistency models, we propose to operate on multiple stages, which simplifies the modelling task and improves performance significantly. On the other hand, TRACT limits itself to distillation and uses the self-evaluation from consistency models to distill models over multiple stages. The stages are progressively reduced to either one or two stages and thus steps. The end-goal of TRACT is again to sample in either one or two steps, whereas we believe better results can be obtained by optimizing for a slightly larger number of steps. We show that this more conservative target, in combination with our improved sampler and annealed schedule, leads to significant improvements in terms of image quality that closes the gap between sample quality of standard diffusion and low-step diffusion-inspired approaches.
Earlier, DDIM (Song et al., 2021a) showed that deterministic samplers degrade more gracefully than the stochastic sampler used by Ho et al. (2020) when limiting the number of sampling steps. Karras et al. (2022) proposed a second order Heun sampler to reduce the number of steps (and function evaluations), while Jolicoeur-Martineau et al. (2021) studied different SDE integrators to reduce function evaluations. Zheng et al. (2023) use specialized architectures to distill the ODE trajectory from a pre-created noise-sample pair dataset. Progressive Distillation (Salimans & Ho, 2022; Meng et al., 2022) distills diffusion models in stages, which limits the number of model evaluations during training while exponentially reducing the required number of sampling steps with the number stages. Luo et al. (2023) distill the knowledge from the diffusion model into a single-step model.
Other methods inspired by diffusion such as Rectified Flows (Liu et al., 2023) and Flow Matching (Lipman et al., 2023) have also tried to reduce sampling times. In practice however, flow matching and rectified flows are generally used to map to a standard normal distribution and reduce to standard diffusion. As a consequence, on its own they still require many evaluation steps. In Rectified Flows, a distillation approach is proposed that does reduce sampling steps more significantly, but this comes at the expense of sample quality.
@@
ImageNet64 ImageNet128
Steps Train Distill Train Distill
Base Consistency Model 1 7.2 4.3 16.0 8.5
MultiStep CM (ours) 2 2.7 2.0 6.0 3.1
MultiStep CM (ours) 4 1.8 1.7 4.0 2.4
MultiStep CM (ours) 8 1.5 1.6 3.3 2.1
MultiStep CM (ours) 16 1.5 1.5 3.4 2.0
Diffusion (aDDIM) 512 1.5 2.2
Large Consistency Model 1 6.4 3.2 14.5 7.0
MultiStep CM (ours) 2 2.3 1.9 4.2 3.1
MultiStep CM (ours) 4 1.6 1.6 2.7 2.3
MultiStep CM (ours) 8 1.5 1.4 2.2 2.1
MultiStep CM (ours) 16 1.6 1.4 2.3 2.0
Diffusion (aDDIM) 512 1.4 2.2
Table 2: Ablation of CD on Image128 with and without annealing the steps on ImageNet128.
@@
Table 3: Comparison between Progressive Distillation (Salimans & Ho, 2022) and CT and CD on ImageNet64 on the base model variant. Reimplementation of PD with modern optimizations.
@@ 5.1 Quantitative Evaluation on ImageNet
Further we are ablating whether annealing the step schedule is important to attain good performance. As can be seen in Table 2, it is especially important for low multistep models to anneal the schedule. In these experiments, annealing always achieves better performance than tests with constant steps at $128,256,1024$. As more multisteps are taken, the importance of the annealing schedule is less important.
Figure 4: Yet another qualititative comparison between a multistep consistency and diffusion model. Top: ours, this time using a standard DDIM distilled 16-step concistency model (3.2 secs). Bottom: generated samples using a 100-step DDIM diffusion model (39 secs). Both models use the same initial noise.
Table 4: Literature Comparison on ImageNet64 and 128.
Method NFE FID
@@ 5.2 Qualitative Evaluation on Text to Image modelling
In addition to the quantitative analysis on ImageNet, we study the effects on a text-to-image model by directly comparing samples. We first train a 20B parameter diffusion model on text-to-image pairs with a T5 XXL paper following (Saharia et al., 2022) for 1.3 million steps. Then, we distill a 16-step consistency model using the DDIM sampler. In Figure 2 and 3 we compare samples from our 16-step CD aDDIM distilled model to the original 100-step DDIM sampler. Because the random seed is shared we can easily compare the samples between these models, and we can see that there are generally minor differences. In our own experience, we often find certain details more precise, at a slight cost of overall construction. Another comparison in Figure 4 shows the difference between a DDIM distilled model (equivalent to $\eta=0$ in aDDIM) and the standard DDIM sampler. Again we see many similarities when sharing the same initial random seed.

### v2

@@ Abstract
In this paper we propose Multistep Consistency Models: A unification between Consistency Models (Song et al., 2023) and TRACT (Berthelot et al., 2023) that can interpolate between a consistency model and a diffusion model: a trade-off between sampling speed and sampling quality. Specifically, a 1-step consistency model is a conventional consistency model whereas a $\infty$-step consistency model is a diffusion model.
Multistep Consistency Models work really well in practice. By increasing the sample budget from a single step to 2-8 steps, we can train models more easily that generate higher quality samples, while retaining much of the sampling speed benefits. Notable results are 1.4 FID on Imagenet 64 in 8 step and 2.1 FID on Imagenet128 in 8 steps with consistency distillation, using simple losses without adversarial training. We also show that our method scales to a text-to-image diffusion model, generating samples that are close to the quality of the original model.
@@ 2 Background: Diffusion Models
Diffusion models are specified by a destruction process that adds noise to destroy data: ${\bm{z}}_{t}=\alpha_{t}{\bm{x}}+\sigma_{t}{\bm{{\epsilon}}}_{t}$ where ${\bm{{\epsilon}}}_{t}\sim\mathcal{N}(0,1)$. Typically for $t\to 1$, ${\bm{z}}_{t}$ is approximately distributed as a standard normal and for $t\to 0$ it is approximately ${\bm{x}}$. In terms of distributions one can write the diffusion process as: $q({\bm{z}}_{t}|{\bm{x}})=\mathcal{N}({\bm{z}}_{t}|\alpha_{t}{\bm{x}},\sigma_{t})$.
Following (Sohl-Dickstein et al., 2015; Ho et al., 2020) we will let $\sigma_{t}^{2}=1-\alpha_{t}^{2}$ (variance preserving). As shown in Kingma et al. (2021), the specific values of $\sigma_{t}$ and $\alpha_{t}$ do not really matter. Whether the process is variance preserving or exploding or something else, they can always be re-parameterized into the other form. Instead, it is their ratio that matters and thus it can be helpful to define the signal-to-noise ratio, i.e. $\mathrm{SNR}(t)=\alpha_{t}^{2}/\sigma_{t}^{2}$. To sample from these models, one uses the denoising equation:
$\small q({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{x}})=\mathcal{N}({\bm{z}}_{s}|\mu_{t\to s}({\bm{z}}_{t},{\bm{x}}),\sigma_{t\to s})$ (1)
@@
$\small||f({\bm{z}}_{t},t)-\operatorname{nograd}(f(\operatorname{DDIM}_{t\to s}({\bm{x}}_{\mathrm{teacher}},{\bm{z}}_{t}),s))||,$ (3)
@@
$\small{\bm{z}}_{s}=\operatorname{DDIM}_{t\to s}({\bm{x}},{\bm{z}}_{t})=\alpha_{s}{\bm{x}}+(\sigma_{s}/\sigma_{t})({\bm{z}}_{t}-\alpha_{t}{\bm{x}})$ (4)
@@
$\small{\bm{x}}=\operatorname{invDDIM}_{t\to s}({\bm{z}}_{s},{\bm{z}}_{t})=\frac{{\bm{z}}_{s}-\frac{\sigma_{s}}{\sigma_{t}}{\bm{z}}_{t}}{\alpha_{s}-\alpha_{t}\frac{\sigma_{s}}{\sigma_{t}}}.$ (5)
@@ 3 Multistep Consistency Models
Algorithm 1 Training Multistep CMs
@@ 3.1 General description
$\small||\operatorname{DDIM}_{t\to t_{\mathrm{step}}}(f({\bm{z}}_{t},t),{\bm{z}}_{t})-\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}}||,$ (6)
where $\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}}=\operatorname{DDIM}_{s\to t_{\mathrm{step}}}(\operatorname{nograd}f({\bm{z}}_{s},s))$ and where the teaching step ${\bm{z}}_{s}=\operatorname{aDDIM}_{t\to s}(x,{\bm{z}}_{t})$. For now it suffices to think $\operatorname{aDDIM}$ as $\operatorname{DDIM}$. It will be described in detail in section 3.2. In fact, one can drop-in any deterministic sampler (or integrator) in place of $\operatorname{aDDIM}$ in the case of distillation.
@@
$\small||\hat{{\bm{x}}}_{\mathrm{diff}}||=||f({\bm{z}}_{t},t)-\operatorname{invDDIM}_{t\to t_{\mathrm{step}}}(\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}},{\bm{z}}_{t})||.$ (7)
@@
$\small||f({\bm{z}}_{t},t)-\operatorname{invDDIM}_{t_{\mathrm{step}}\to t}({\bm{z}}_{t_{\mathrm{step}}})||.$ (8)
@@
$w(t)||{\bm{x}}_{\mathrm{diff}}||=\operatorname{invDDIM}_{t\to s}(\hat{{\bm{z}}}_{s},{\bm{z}}_{t})-\hat{{\bm{x}}}=w(t)||{\bm{x}}-\hat{{\bm{x}}}||$ (9)
@@
Algorithm 2 Sampling from Multistep CMs
  Sample ${\bm{z}}_{T}\sim\mathcal{N}(0,{\mathbf{I}})$
     ${\bm{z}}_{s}=\operatorname{DDIM}_{t\to s}(f({\bm{z}}_{t},t),{\bm{z}}_{t})$
  end for
@@
$\small{\bm{z}}^{*}_{s}=\alpha_{s}{\bm{x}}^{*}+\sigma_{s}\frac{{\bm{z}}_{t}-\alpha_{t}{\bm{x}}^{*}}{\sigma_{t}}=(\alpha_{s}-\frac{\alpha_{t}\sigma_{s}}{\sigma_{t}}){\bm{x}}^{*}+\frac{\sigma_{s}}{\sigma_{t}}{\bm{z}}_{t}.$ (10)
@@
$\small{\bm{z}}^{\text{DDIM}}_{s}=(\alpha_{s}-\alpha_{t}\sigma_{s}/\sigma_{t})\hat{{\bm{x}}}+(\sigma_{s}/\sigma_{t}){\bm{z}}_{t},$ (11)
@@
$\small||{\bm{z}}^{*}_{s}||^{2}-||{\bm{z}}^{\text{DDIM}}_{s}||^{2}=\text{trace}(\mathrm{Var}[{\bm{z}}_{s}|{\bm{z}}_{t}]),$ (12)
@@
$\small\mathrm{Var}[{\bm{z}}_{s}|{\bm{z}}_{t}]=(\alpha_{s}-\alpha_{t}\sigma_{s}/\sigma_{t})^{2}\cdot\mathrm{Var}[{\bm{x}}|{\bm{z}}_{t}],$ (13)
@@
Figure 3: Comparison of the vanilla and adjusted DDIM sampler for the small ImageNet 128 diffusion model without distillation.
Currently, the best sample quality is achieved with stochastic samplers, which can be tuned to add exactly enough noise to undo the oversmoothing caused by numerical integration. However, current distillation methods are not well suited to distilling these stochastic samplers directly. Here we therefore propose a new deterministic sampler that aims to achieve the norm increasing effect of noise addition in a deterministic way. It turns out we can do this by making a simple adjustment to the DDIM sampler, and we therefore call our new method Adjusted DDIM (aDDIM). Our modification is heuristic and is not more theoretically justified than the original DDIM sampler. However, empirically we find aDDIM to work very well leading to improved FID scores (Fig. 3) and thus a stronger deterministic teacher.
@@
$\small{\bm{z}}^{\text{aDDIM}}_{s}=\alpha_{s}\hat{{\bm{x}}}+\sqrt{\sigma_{s}^{2}+\text{tr}(\mathrm{Var}[{\bm{z}}_{s}|{\bm{z}}_{t}])/||\hat{{\bm{{\epsilon}}}}||^{2}}\cdot\hat{{\bm{{\epsilon}}}}.$ (14)
@@
Figure 4: Another qualititative comparison between a multistep consistency and teacher, using the same prompt. Top: ours, a distilled 16-step concistency model (3.2 secs). Bottom: generated samples using a 100-step DDIM diffusion model (39 secs). Both models use the same initial noise.
@@ 4 Related Work
Multistep Consistency Models are a direct combination of (Song et al., 2023; Song & Dhariwal, 2023) and TRACT (Berthelot et al., 2023). Compared to consistency models, we propose to operate on multiple stages, which simplifies the modelling task and improves performance significantly. On the other hand, TRACT limits itself to distillation and uses the self-evaluation from consistency models to distill models over multiple stages. The stages are progressively reduced to either one or two stages and thus steps. The end-goal of TRACT is again to sample in either one or two steps, whereas we believe better results can be obtained by optimizing for a slightly larger number of steps. We show that this more conservative target, in combination with our improved sampler and annealed schedule, leads to significant improvements in terms of image quality that closes the gap between sample quality of standard diffusion and low-step diffusion-inspired approaches.
Earlier, DDIM (Song et al., 2021a) showed that deterministic samplers degrade more gracefully than the stochastic sampler used by Ho et al. (2020) when limiting the number of sampling steps. Karras et al. (2022) proposed a second order Heun sampler to reduce the number of steps (and function evaluations), while Jolicoeur-Martineau et al. (2021) studied different SDE integrators to reduce function evaluations. Progressive Distillation (Salimans & Ho, 2022; Meng et al., 2022) distills diffusion models in stages, which limits the number of model evaluations during training while exponentially reducing the required number of sampling steps with the number stages.
Other methods inspired by diffusion such as Rectified Flows (Liu et al., 2023a) and Flow Matching (Lipman et al., 2023) have also tried to reduce sampling times. In practice however, flow matching and rectified flows are generally used to map to a standard normal distribution and reduce to standard diffusion. As a consequence, on its own they still require many evaluation steps. In Rectified Flows, a distillation approach is proposed that does reduce sampling steps more significantly, but this comes at the expense of sample quality.
Adversarial distillation
Distillation to a few steps was very difficult to do using only simple distance metrics. Therefore, many works resort to a form of adversarial training. For example Luo et al. (2023) distill the knowledge from the diffusion model into a single-step model and Zheng et al. (2023) use specialized architectures to distill the ODE trajectory from a pre-created noise-sample pair dataset. A very similar approach to ours is Consistency Trajectory Models (CTMs) (Kim et al., 2023), which are trained to arbitrarily integrate to a given timestep. This is implemented by modifying the inputs of the denoising network to include an endpoint of the integration. Although CTMs produce very high quality image samples in few steps, their performance relies on adversarial training: Without it, CTMs cannot produce great samples and have a considerable gap in FID score. In contrast, our Multistep CMs can be trained with simple distance metrics and still achieve very good FID scores under few sampling steps. A possible explanation is that it is much easier to learn a handful of fixed integration trajectories (Multistep CMs) instead of every possible integration with arbitrary endpoints (CTMs). Another advantage of Multistep CMs is that the inputs to the denoising network are not changed, making fine-tuning of existing diffusion models to Multistep CMs very straightforward.
@@
Small Large
ImageNet64 ImageNet128 ImageNet64 ImageNet128
Steps Train Distill Train Distill Train Distill
Base Consistency Model 1 7.2 4.3 16.0 8.5 6.4 3.2 14.5 7.0
MultiStep CM (ours) 2 2.7 2.0 6.0 3.1 2.3 1.9 4.2 3.1
MultiStep CM (ours) 4 1.8 1.7 4.0 2.4 1.6 1.6 2.7 2.3
MultiStep CM (ours) 8 1.5 1.6 3.3 2.1 1.5 1.4 2.2 2.1
MultiStep CM (ours) 16 1.5 1.5 3.4 2.0 1.6 1.4 2.3 2.0
Diffusion (aDDIM) 512 1.5 2.2 1.4 2.2
Table 2: Ablation of CD on Image128 with and without annealing the teacher steps on ImageNet128. Annealing the teacher stepsize improves the performance.
@@
Table 3: Comparison between PD (Salimans & Ho, 2022) and CT/CD on ImageNet64 on the small model.
@@ 5.1 Quantitative Evaluation on ImageNet
Table 4: Literature Comparison on ImageNet.
Method NFE FID non-adv
@@
Further we ablate whether annealing the step schedule is important to attain good performance. As can be seen in Table 2, it is especially important for low multistep models to anneal the schedule. In these experiments, annealing always achieves better performance than tests with constant steps at $128,256,1024$. As more multisteps are taken, the importance of the annealing schedule decreases.
@@ 5.2 Qualitative Evaluation on Text to Image modelling
In addition to the analysis on ImageNet, we study the effects on a text-to-image models. We distill a 16-step consistency model a base teacher model. In Table 5 one can see that Multistep CD is able to distill its teacher almost perfectly in terms of FID. The loss of clip score can be attributed to the guidance distillation, which a baseline 256-step student model also has trouble distilling. Compared to the guidance-distilled baseline, the 16-CD model has no loss in performance measured in CLIP and FID on the low guidance setting (and for the high guidance setting only a minor degradation in FID). Even the 8-step CD model attains an impressive FID score of 8.1, which is well below the existing literature.
In Figure 2 and 5 we compare samples from our 16-step CD aDDIM distilled model to the original 100-step DDIM sampler. Because the random seed is shared we can easily compare the samples between these models, and we can see that there are generally minor differences. In our own experience, we often find certain details more precise, at a slight cost of overall construction. Another comparison in Figure 4 shows the difference between a DDIM distilled model (equivalent to $\eta=0$ in aDDIM) and the standard DDIM sampler. Again we see many similarities when sharing the same initial random seed.
Table 5: Text to Image performance. Note that when 8/16-step Consistency is compared to a teacher model that is only guidance distilled at 256 steps, there is practically no performance loss.
[…]
SDv1.5 (Rombach et al., 2022) low g (from DMD) 512 8.8 -
                                         high g (from DMD) 512 13.5 0.322
DMD (low guidance) (Yin et al., 2023) 1 11.5 -
      (high guidance) 1 14.9 0.32
UFOGen (Xu et al., 2023) 1 12.8 22.5 0.311
4 22.1 0.307
InstaFlow-1.7B (Liu et al., 2023b) 1 11.8 22.4 0.309
PeRFlow (Yan et al., 2024) 4 11.3
Teacher Diffusion Model g=0.5 (ddpm) 256 7.9 13.6 0.305
     guidance distilled (ddim) 256 8.2 13.8 0.300
Multistep-CD (teacher g=0.5) 4 8.7 14.4 0.298
8 8.1 13.8 0.300
16 7.9 13.9 0.300
Teacher Diffusion Model g=3 (ddpm) 256 12.7 18.1 0.315
     guidance distilled (ddim) 256 13.9 19.0 0.312
Multistep-CD (teacher g=3) 4 12.4 18.1 0.311
8 13.9 19.6 0.311
16 14.4 20.0 0.312